# DiaBlo: Diagonal Blocks Are Sufficient For Finetuning

### v1

@@ Abstract
Fine-tuning is a critical step for adapting large language models (LLMs) to domain-specific downstream tasks. To mitigate the substantial computational and memory costs of full-model fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to update only a small subset of model parameters. However, performance gaps between PEFT approaches and full-model fine-tuning still exist. In this work, we present DiaBlo, a simple yet effective PEFT approach that updates only the diagonal blocks of selected model weight matrices. Unlike Low-Rank Adaptation (LoRA) and its variants, DiaBlo eliminates the need for low-rank matrix products, thereby avoiding the reliance on auxiliary initialization schemes or customized optimization strategies to improve convergence. This design leads to stable and robust convergence while maintaining comparable memory efficiency and training speed to LoRA. We conduct extensive experiments across a range of tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—to evaluate the effectiveness and efficiency of DiaBlo. Across these benchmarks, DiaBlo demonstrates strong and consistent performance while maintaining high memory efficiency and fast fine-tuning speed. Codes are available at https://github.com/ziyangjoy/DiaBlo.
@@ 1 Introduction
Large language models (LLMs) Achiam2023GPT4TR ; Brown2020LanguageMA ; Touvron2023LLaMAOA have achieved remarkable success across a wide range of natural language processing (NLP) tasks, including reasoning, generation, and alignment. These models, often consisting of billions of parameters, are typically pre-trained on broad, general-purpose corpora. However, to adapt them to specific downstream tasks or domains, fine-tuning is essential. Full fine-tuning, which updates all parameters of the model, has been shown to yield strong performance but is prohibitively expensive in terms of computational cost, memory usage, and storage—especially on resource-constrained devices.
To address these challenges, Parameter-Efficient Fine-Tuning (PEFT) methods Ding2023ParameterefficientFO ; Han2024ParameterEfficientFF have emerged as a promising alternative. These methods aim to retain the performance benefits of full fine-tuning while updating only a small fraction of the model’s parameters. Early approaches, such as Prompt Tuning Lester2021ThePO and Prefix Tuning Li2021PrefixTuningOC , introduce task-specific embeddings or continuous prompts that steer model behavior without modifying its original weights. These methods are lightweight and straightforward to implement, making them suitable for low-resource adaptation scenarios. However, they often struggle with scalability and limited expressiveness, especially on complex tasks. To improve performance, more structured methods like Low-Rank Adaptation (LoRA) Hayou2024LoRAEL ; Hu2021LoRALA have been proposed. LoRA adds trainable low-rank matrices to the pretrained weight, significantly reducing the number of trainable parameters. It has demonstrated strong performance across a variety of applications. Despite its efficiency, LoRA and its variants often encounter issues such as unstable convergence and performance degradation. A growing number of work has proposed various extensions to LoRA, including tailored initialization schemes Meng2024PiSSAPS ; Wang2024MiLoRAHM ; Wang2024LoRAGALA and structure-specific optimization strategies DoRA ; Wang2024LoRAProAL . Nonetheless, such improvements may lead to greater algorithmic complexity, potentially impacting implementation and efficiency.
@@
In this work, we propose DiaBlo, a simple yet effective PEFT framework that fine-tunes only the diagonal blocks of the model’s weight matrices. Unlike LoRA, which introduces low-rank adaptation through the product of two trainable matrices—a structure that often requires careful initialization and customized optimization to ensure stable training—DiaBlo directly updates a structured subset of the model’s original weight parameters. By avoiding the use of matrix products, DiaBlo eliminates the inherent optimization difficulties associated with low-rank decomposition. This leads to greater training stability and more reliable convergence without the need for special tricks or tuning. Despite its simplicity, DiaBlo retains the memory and computational efficiency of LoRA while offering improved representational flexibility. It integrates easily into standard training pipelines and consistently outperforms existing PEFT methods across a broad range of tasks, demonstrating that selectively updating diagonal blocks is a powerful and practical alternative for parameter-efficient fine-tuning. Figure 1 illustrates the difference between full fine-tuning, LoRA, and DiaBlo.
Geometric interpretation of DiaBlo is based on the observation that LoRA’s representational capacity is inherently directionally and globally structured. By introducing a diagonal matrix into the PEFT model, we can apply axis-aligned, per-dimension scaling that LoRA cannot express. Geometrically, this means that while LoRA adjusts the weight matrix by steering it within a constrained low-rank plane, the diagonal component expands this plane into a sheared, anisotropically rescaled region, enabling the model to fine-tune the sensitivity of individual features. This additional flexibility is especially valuable in transfer learning for reasoning tasks, where subtle, independent shifts in feature activation can significantly improve generalization with minimal parameter overhead. Thus, the diagonal augmentation enhances the expressiveness while maintaining high efficiency.
We evaluate DiaBlo across a range of comprehensive tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—to rigorously assess its effectiveness and efficiency. Across all benchmarks, DiaBlo consistently achieves state-of-the-art performance among PEFT approaches, often narrowing or even surpassing the gap to full fine-tuning. These results highlight not only the expressive power of diagonal block updates but also the method’s robustness across diverse domains. DiaBlo demonstrates strong convergence behavior, efficient resource usage, and broad applicability, making it a compelling alternative to existing parameter-efficient methods.
@@
  • A novel PEFT framework: We propose DiaBlo, a simple yet effective method that fine-tunes only the diagonal blocks of model weight matrices. This structured approach enables efficient adaptation while achieving strong performance across a variety of downstream tasks.
  • Stable optimization behavior: DiaBlo directly updates diagonal blocks using standard training pipelines, without relying on matrix product structures or specialized initialization, resulting in more stable and reliable fine-tuning.
  • Strong empirical performance: Through extensive experiments on tasks including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment, we demonstrate that DiaBlo consistently achieves strong downstream performance, often surpassing LoRA and its variants in accuracy.
  • Memory and computational efficiency: DiaBlo retains the low memory footprint and fast training speed of LoRA and other PEFT methods, ensuring practical usage for large-scale fine-tuning.
@@ 2 Background and Related Works
Early PEFT techniques, such as Prompt Tuning Han2021PTRPT ; Lester2021ThePO and Prefix Tuning Li2021PrefixTuningOC , introduce learnable embeddings or token sequences that steer the model’s behavior without modifying the core parameters. These methods are lightweight and easy to integrate, offering low computational overhead. While effective in low-resource or few-shot scenarios, prompt-based methods often struggle to match the performance of full fine-tuning on more complex or general-purpose tasks due to their limited expressiveness and adaptability. Adapter-based approaches Houlsby2019ParameterEfficientTL ; Rckl2020AdapterDropOT ; Wang2020KAdapterIK insert small, trainable bottleneck layers into the transformer architecture. These modules enable task-specific learning while keeping most of the model’s parameters frozen. Variants of adapter tuning further improve performance through mechanisms such as attention fusion or parameter sharing. Although adapters may reach a good performance, they often require changes to model architecture and introduce additional inference-time latency. In addition, there exist matrix or tensor decomposition-based approaches Lingam2024SVFTPF ; Liu2025CoLACP ; Wang2024SVDLLMTS ; Yang2024LoRETTALE ; Yang2024CoMERACA ; Yang2023QuantizationAwareAT to reduce training or finetuning costs.
Low-Rank Adaptation and Variants
@@
Low-Rank Adaptation (LoRA) Hu2021LoRALA has emerged as one of the most popular PEFT methods. LoRA injects low-rank matrices into the update paths of existing model weights, significantly reducing the number of trainable parameters. Its effectiveness and simplicity have led to wide adoption, but its core reliance on the product of two low-rank matrices introduces optimization challenges. This structure can restrict representational flexibility and make training sensitive to initialization and optimization algorithms.
@@
Several recent variants have been proposed to enhance the performance of LoRA. DoRA DoRA introduces a novel optimization strategy that decouples the magnitude and direction of low-rank updates, leading to more stable training dynamics and better performance. Pissa Meng2024PiSSAPS and MiLoRA Wang2024MiLoRAHM focus on different initialization techniques for low-rank adapters. Pissa initializes the low-rank matrices using the largest singular values and singular vectors of weights, thereby preserving the most significant components of the weight matrix for faster convergence. On the other hand, MiLoRA takes a different approach by utilizing the smallest singular values during initialization. This strategy captures the finer details of the weight matrix. Additionally, LoRA-GA Wang2024LoRAGALA proposes to initialize low-rank matrices to align with gradients of full fine-tuning at the first step. While these LoRA-based methods have achieved notable improvements, they often introduce additional complexity in design and implementation. Furthermore, the dependence on matrix-product-based updates continues to present challenges for stable and efficient optimization. There exist many other variants of LoRA. We refer readers to Kalajdzievski2023ARS ; Valipour2022DyLoRAPT ; Wang2024RoseLoRARA ; Zhang2023AdaLoRAAB ; Zhong2024NEATNP . QLoRA dettmers2023qlora reduces the memory footprint of fine-tuning by applying low-rank adaptation on quantized models. Building on this idea, subsequent works GPTQLoRA ; deng2025cloq ; li2023loftq ; liao2024apiq design customized quantization algorithms and develop effective initialization strategies for the low-rank adaptation components, aiming to improve fine-tuning performance.
@@ 3.1 DiaBlo: Finetuning Diagonal Blocks
Despite the success of LoRA-based fine-tuning methods, low-rank factorizations remain difficult to optimize in practice. In computational mathematics, problems such as low-rank approximation and low-rank matrix completion are well known to be non-convex and challenging Ma2019BeyondPB ; Tong2020AcceleratingIL ; Tu2015LowrankSO . These difficulties are reflected in LoRA-based approaches, where the product of two trainable low-rank matrices can lead to unstable gradient flow and affect convergence. As a result, much recent work has focused on designing better initialization schemes and custom optimization strategies to stabilize training. The development of DiaBlo is motivated by these limitations. To this end, DiaBlo directly fine-tunes only the diagonal blocks of model weight matrices, avoiding matrix-product parameterizations entirely. This design simplifies the optimization landscape, improves training stability, and enables effective adaptation without introducing additional complexity.
@@
where $\mathbf{W}_{ij}\in\mathbb{R}^{d_{1}\times d_{2}},d_{1}=\frac{m_{1}}{N},d_{2}=\frac{m_{2}}{N}$. During finetuning, only the diagonal blocks of $\mathbf{W}$, i.e., $\mathbf{W}_{11},\ldots,\mathbf{W}_{NN}$, are trainable while all other blocks $\{\mathbf{W}_{ij}\}_{i\neq j}$ are frozen. DiaBlo also works when $N$ is not a common factor of $m_{1},m_{2}$. In that case, we may expand the pad the weight $\mathbf{W}$ into a proper size and then select the corresponding diagonal blocks.
@@
$\mathbf{Y}=\mathbf{X}\mathbf{W}+\mathbf{X}\mathbf{D}.$ (1)
@@
[…]
@@
LoRA and its variants typically represent adaptation as a low-rank product $\mathbf{A}\mathbf{B}$. Standard initialization schemes set $\mathbf{A}$ using Kaiming uniform and $\mathbf{B}$ as zero, ensuring $\mathbf{A}\mathbf{B}=0$ while the gradients to $\mathbf{A},\mathbf{B}$ do not vanish. However, such initialization often leads to suboptimal performance, as observed in recent studies Meng2024PiSSAPS ; Wang2024MiLoRAHM ; Wang2024LoRAGALA . Consequently, variants like Pissa Meng2024PiSSAPS , MiLoRA Wang2024MiLoRAHM , and LoRA-GA Wang2024LoRAGALA propose alternative initialization techniques to improve optimization behavior—e.g., Pissa uses large singular values, while MiLoRA leverages small singular values for initialization. In contrast, DiaBlo does not involve matrix products and can be initialized more straightforwardly. We simply initialize $\mathbf{\mathcal{D}}$ as an all-zero tensor. This avoids issues related to vanishing gradients or entangled parameter updates, enabling stable and efficient training without the need for specialized initialization and optimization techniques.
3.2 Comparison with LoRA
Gradient Analysis

### v2

@@ Abstract
Fine-tuning is a critical step for adapting large language models (LLMs) to domain-specific downstream tasks. To mitigate the substantial computational and memory costs of full-model fine-tuning, Parameter-Efficient Fine-Tuning (PEFT) methods have been proposed to update only a small subset of model parameters. However, performance gaps between PEFT approaches and full-model fine-tuning still exist. In this work, we present DiaBlo, a simple yet effective PEFT approach that updates only the diagonal blocks of selected model weight matrices. Unlike Low-Rank Adaptation (LoRA) and its variants, DiaBlo eliminates the need for low-rank matrix products, thereby avoiding the reliance on auxiliary initialization schemes or customized optimization strategies to improve convergence. This design leads to stable and robust convergence while maintaining comparable memory efficiency and training speed to LoRA. Moreover, we provide theoretical guarantees showing that, under mild low-rank conditions, DiaBlo is more expressive than LoRA in the linear problem and converges to a stationary point of the general nonlinear full fine-tuning. Through extensive experiments across a range of tasks—including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment—we show that fine-tuning only diagonal blocks is sufficient for strong and consistent performance. DiaBlo not only achieves competitive accuracy but also preserves high memory efficiency and fast fine-tuning speed. Codes are available at https://github.com/ziyangjoy/DiaBlo.
footnotetext: * Equal contribution   Corresponding author
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Brown et al., 2020; Touvron et al., 2023a) have achieved remarkable success across a wide range of natural language processing tasks, including reasoning, generation, and alignment. These models, often consisting of billions of parameters, are typically pre-trained on broad, general-purpose corpora. However, to adapt them to specific downstream tasks or domains, fine-tuning is essential. Full fine-tuning, which updates all parameters of the model, has been shown to yield strong performance but is prohibitively expensive in terms of computational cost, memory usage, and storage—especially on resource-constrained devices.
To address these challenges, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Han et al., 2024) have emerged as a promising alternative. These methods aim to retain the performance benefits of full fine-tuning while updating only a small fraction of the model’s parameters. Early approaches, such as Prompt Tuning (Lester et al., 2021) and Prefix Tuning (Li and Liang, 2021), introduce task-specific embeddings or continuous prompts that steer model behavior without modifying its original weights. These methods are lightweight and straightforward to implement, making them suitable for low-resource adaptation scenarios. However, they often struggle with scalability and limited expressiveness, especially on complex tasks. To improve performance, more structured methods like Low-Rank Adaptation (LoRA) (Hayou et al., 2024; Hu et al., 2021) have been proposed. LoRA adds trainable low-rank matrices to the pretrained weight, significantly reducing the number of trainable parameters. It has demonstrated strong performance across a variety of applications. Despite its efficiency, LoRA and its variants often encounter issues such as unstable convergence and performance degradation. A growing number of work has proposed various extensions to LoRA, including tailored initialization schemes (Meng et al., 2024; Wang et al., 2024a; d) and structure-specific optimization strategies (Liu et al., 2024; Wang and Liang, 2024). Nonetheless, such improvements may lead to greater algorithmic complexity, potentially impacting implementation and efficiency. Sparsity-based methods, in contrast, aim to fine-tune only a subset of weight entries. Yet, most existing approaches rely on unstructured sparsity—through either random masking (Deng et al., 2024; Rios et al., 2025; Xu and Zhang, 2024) or importance-based selection (He et al., 2024)—which not only increases time complexity but also results in patterns that are difficult to exploit efficiently on modern hardware. $\text{S}^{2}$FT (Yang et al., 2024a) is an efficient LLM adaptation framework that selectively updates sparse components while computing on dense submatrices. SparseLoRA (Khaki et al., 2025) employs a lightweight SVD-based estimator to identify sparse weights for loss and gradient computation, thereby accelerating LLM fine-tuning.
@@
In this work, we propose DiaBlo, a simple yet effective PEFT framework that fine-tunes only the diagonal blocks of the model’s weight matrices. Figure 1 illustrates the difference between full fine-tuning, LoRA, and DiaBlo. Unlike LoRA, which introduces low-rank adaptation through the product of two trainable matrices that often requires careful initialization and customized optimization to ensure stable training, DiaBlo directly updates a structured subset of the model’s original weight parameters. By avoiding the use of matrix products, DiaBlo eliminates the inherent optimization difficulties associated with low-rank decomposition. This leads to greater training stability and more reliable convergence without the need for special tricks or tuning. Unlike existing fine-tuning approaches based on unstructured sparsity, DiaBlo’s structured diagonal blocks preserve memory and computational efficiency while providing hardware-friendly patterns and enhanced representational flexibility. It integrates easily into standard training pipelines and consistently outperforms existing PEFT methods across a broad range of tasks, demonstrating that selectively updating diagonal blocks is a powerful and practical alternative for parameter-efficient fine-tuning.
As shown in Figure 2, DiaBlo clearly outperforms other PEFT baselines in both commonsense and arithmetic reasoning with the full-precision LLaMA2-7B model, and it maintains this advantage under quantized settings. Notably, DiaBlo demonstrates higher stability in 4-bit and 2-bit arithmetic reasoning tasks, where competing methods show significant performance degradation. These results highlight not only the expressive power of diagonal block updates but also the method’s robustness across diverse domains. DiaBlo demonstrates strong convergence behavior, efficient resource usage, and broad applicability, making it a compelling alternative to existing parameter-efficient methods.
@@
  • Sufficiency of Diagonal Block: We propose DiaBlo, which fine-tunes only the diagonal blocks of model weight matrices. Through extensive experiments on tasks including commonsense reasoning, arithmetic reasoning, code generation, and safety alignment, we demonstrate that DiaBlo consistently achieves strong downstream performance, often surpassing LoRA and its variants in accuracy.
  • Theoretical Guarantees: We provide theoretical justifications for DiaBlo. We show that, under mild low-rank assumptions on activations and gradients that often approximately hold in practice, DiaBlo converges to a stationary solution of full fine tuning (FT), provided the diagonal block size is sufficiently large. In the linear least squares setting, we further prove that DiaBlo converges to a global minimizer of the full FT objective and is strictly more expressive than LoRA under the same parameter budget.
  • Simple Optimization: DiaBlo directly updates diagonal blocks using standard training pipelines, without relying on matrix product structures or specialized initialization, resulting in more stable and reliable fine-tuning.
  • High computation efficiency: DiaBlo retains the low memory footprint and fast training of vanilla LoRA and other PEFT methods, ensuring practical usage for large fine-tuning.
(a) Finetuning full-precision LLaMA2-7B
(b) Finetuning quantized LLaMA2-7B
Figure 2: Comparison of DiaBlo with other PEFT methods on finetuning LLaMA2-7B.
@@ 2 Background and Related Works
Early PEFT techniques, such as Prompt Tuning (Han et al., 2021; Lester et al., 2021) and Prefix Tuning (Li and Liang, 2021), introduce learnable embeddings or token sequences that steer the model’s behavior without modifying the core parameters. These methods are lightweight and easy to integrate, offering low computational overhead. While effective in low-resource or few-shot scenarios, prompt-based methods often struggle to match the performance of full fine-tuning on more complex tasks due to their limited expressiveness and adaptability. Adapter-based approaches (Houlsby et al., 2019; Rücklé et al., 2020; Wang et al., 2020) insert small, trainable bottleneck layers into the transformer architecture. These modules enable task-specific learning while keeping most of the model’s parameters frozen. Variants of adapter tuning further improve performance through mechanisms such as attention fusion or parameter sharing. Although adapters may reach a good performance, they often require changes to model architecture and introduce additional inference-time latency. In addition, there exist decomposition-based approaches (Lingam et al., 2024; Liu et al., 2025; Wang et al., 2024e; Yang et al., 2024b; c; 2023) to reduce training or finetuning costs.
@@
Low-Rank Adaptation (LoRA) (Hu et al., 2021) has emerged as one of the most popular PEFT methods. LoRA injects low-rank matrices into the update paths of existing model weights, significantly reducing the number of trainable parameters. Its effectiveness and simplicity have led to wide adoption, but its core reliance on the product of two low-rank matrices introduces optimization challenges. This structure can restrict representational flexibility and make training sensitive to initialization and optimization algorithms.
@@
Several recent variants have been proposed to enhance the performance of LoRA. DoRA (Liu et al., 2024) introduces a novel optimization strategy that decouples the magnitude and direction of low-rank updates, leading to more stable training dynamics and better performance. Pissa (Meng et al., 2024) and MiLoRA (Wang et al., 2024a) focus on different initialization techniques for low-rank adapters. Pissa initializes the low-rank matrices using the largest singular values and singular vectors of weights, thereby preserving the most significant components of the weight matrix for faster convergence. On the other hand, MiLoRA takes a different approach by utilizing the smallest singular values during initialization. This strategy captures the finer details of the weight matrix. Additionally, LoRA-GA (Wang et al., 2024d) proposes to initialize low-rank matrices to align with gradients of full fine-tuning at the first step. While these LoRA-based methods have achieved notable improvements, they often introduce additional complexity in design and implementation. Furthermore, the dependence on matrix-product-based updates continues to present challenges for stable and efficient optimization. There exist many other variants of LoRA. We refer readers to (Kalajdzievski, 2023; Ponkshe et al., 2024; Valipour et al., 2022; Wang et al., 2024b; Zhang et al., 2023; 2025b; Zhong et al., 2024). QLoRA (Dettmers et al., 2023) reduces the memory footprint of fine-tuning by applying low-rank adaptation on quantized models. Building on this idea, subsequent works (18; Y. Deng, A. Zhang, N. Wang, S. Gurses, Z. Yang, and P. Yin (2025); Y. Li, Y. Yu, C. Liang, P. He, N. Karampatziakis, W. Chen, and T. Zhao (2023); B. Liao, C. Herold, S. Khadivi, and C. Monz (2024)) design customized quantization algorithms and develop effective initialization strategies for the low-rank adaptation components, aiming to improve fine-tuning performance.
@@ 3.1 DiaBlo: Finetuning Diagonal Blocks
Despite the success of LoRA-based fine-tuning methods, low-rank factorizations remain difficult to optimize in practice. In computational mathematics, problems such as low-rank approximation and low-rank matrix completion are well known to be non-convex and challenging (Ma et al., 2019; Tong et al., 2020; Tu et al., 2015). These difficulties are reflected in LoRA-based approaches, where the product of two trainable low-rank matrices can lead to unstable gradient flow and affect convergence. As a result, much recent work has focused on designing better initialization schemes and custom optimization strategies to stabilize training. The development of DiaBlo is motivated by these limitations. To this end, DiaBlo directly fine-tunes only the diagonal blocks of model weight matrices, avoiding matrix-product parameterizations entirely. This design simplifies the optimization landscape, improves training stability, and enables effective adaptation without introducing additional complexity.
@@
where $\mathbf{W}_{ij}\in\mathbb{R}^{d_{1}\times d_{2}},d_{1}=\frac{m_{1}}{N},d_{2}=\frac{m_{2}}{N}$. During finetuning, only the diagonal blocks of $\mathbf{W}$, i.e., $\mathbf{W}_{11},\ldots,\mathbf{W}_{NN}$, are trainable while all other blocks $\{\mathbf{W}_{ij}\}_{i\neq j}$ are frozen. DiaBlo also works when $N$ is not a common factor of $m_{1},m_{2}$. In this case, we set $d_{1}=\left\lceil\frac{m_{1}}{N}\right\rceil,d_{2}=\left\lceil\frac{m_{2}}{N}\right\rceil.$ At inference time, we pad $\mathbf{X}$ with zeros so that its dimension becomes $b\times Nd_{1}$, perform the blockwise computation, and then truncate the resulting output back to the size $b\times m_{2}$. This situation is uncommon in practice because modern LLM architectures purposely choose hidden sizes and intermediate dimensions that are highly composite, often powers of two or products of small primes with powers of two, ensuring many convenient common factors. For instance, the Llama3-8B model has a hidden dimension of $4096=2^{12}$ and an FFN intermediate dimension of $14336=7\times 2^{11}$. This allows us to choose $N=2^{k}$ for $k=1,\ldots,11$. In practice, $N=32,64,128$ are commonly used, as these choices are always common factors and align well with the trainable-parameter budgets of other PEFT methods.
@@
$\mathbf{Y}=\mathbf{X}\mathbf{W}=\mathbf{X}\mathbf{W}_{0}+\mathbf{X}\mathbf{D}.$ (1)
@@
$\texttt{torch.einsum}(bNd_{1},Nd_{1}d_{2}\rightarrow bNd_{2},\mathbf{\mathcal{X}},\mathbf{\mathcal{D}}).$
@@
LoRA and its variants typically represent adaptation as a low-rank product $\mathbf{A}\mathbf{B}$. Standard initialization schemes set $\mathbf{A}$ using Kaiming uniform and $\mathbf{B}$ as zero, ensuring $\mathbf{A}\mathbf{B}=\mathbf{0}$ while the gradients to $\mathbf{A},\mathbf{B}$ do not vanish. However, such initialization often leads to suboptimal performance, as observed in recent studies (Meng et al., 2024; Wang et al., 2024a; d). Consequently, variants like Pissa (Meng et al., 2024), MiLoRA (Wang et al., 2024a), and LoRA-GA (Wang et al., 2024d) propose alternative initialization techniques to improve optimization behavior—e.g., Pissa uses large singular values, while MiLoRA leverages small singular values for initialization. In contrast, DiaBlo does not involve matrix products and can be initialized more straightforwardly. We simply initialize $\mathbf{\mathcal{D}}$ as an all-zero tensor. This avoids issues related to vanishing gradients or entangled parameter updates, enabling stable and efficient training without the need for specialized initialization and optimization techniques.
3.2 Theoretical Analysis
Comparison with LoRA